Convert 41.9degrees and 74.8degrees to the nearest minutes. Please show working
Answer by Alan3354(69443)

Same problem, different numbers.
=========================
Convert 41.9degrees and 74.28degrees to the nearest degree, minutes and seconds
---------------------
0.28 degs = 0.28*60' = 16.8'
0.8' = 0.8*60" = 48"
---> 74degs 16' 48"